# Susan Athey    Emil Palikot

### v4

@@ 6 Counterfactual Program Assignment Policies
The number of spots in each program is limited. The availability of mentors restricts the capacity of Mentoring. In Challenges, we limited the available spots to 300 to accommodate potential operational complexities and logistical problems. Considering the substantial treatment effects of both programs, relaxing these capacity constraints will increase Dare IT’s impact. However, capacity constraints are likely to remain a long-term issue for the Mentoring program and may persist, albeit to a lesser extent, for Challenges. Consequently, Dare IT will continue to face the question of which applicants to accept for each program. Given the heterogeneous treatment effects, admission policies that favor applicants with high predicted benefits could enhance the impact of Dare IT programs. This section aims to quantify these potential gains.
@@
We perform three analyses: (i.) we estimate gains from the introduction of a program admission policy that, given capacity levels in each program, assigns applicants to programs, maximizing the sum of predicted treatment effects of programs’ participants; (ii.) we evaluate the gains from relaxing capacity constraints with the targeting policy in place; (iii.) we compare the overall impact of Mentoring when mentors select the most promising candidates to a case where they select to following alternative strategies.
@@
An assignment policy is a mapping from the characteristics of the applicants $X\in\mathcal{X}$ and the capacity levels of the programs $Q=(Q^{M},Q^{C})\in\mathcal{Q}$ to assignments: $\pi:\mathcal{X}\times\mathcal{Q}\rightarrow\mathcal{P}$, where $\mathcal{P}\equiv\left\{M,C,O\right\}$ and $M$ stands for Mentoring, $C$ for Challenges, and $O$ for Out of Dare IT. $\mathcal{Q}$ is the grid of possible capacity levels under consideration.
Policy estimation and evaluation are performed separately in train and test sets. Across all analyses, we obtain train and test sets by randomly splitting data at the applicant level with equal probability ($\mathcal{I}_{train}$ and $\mathcal{I}_{test}$ are resulting sets of applicants in the train and test data sets, respectively). The objective of policy estimation is to obtain a mapping from applicant characteristics and programs’ capacity levels to program assignments. We use the train set to obtain the estimate of the policy $\pi$, which we denote $\hat{\pi}$.
@@
Next, to evaluate a policy, we proceed in two steps: first, we apply the estimated mapping $\hat{\pi}$ to the test set, where for each $i$ in the test set and $Q\in\mathcal{Q}$, we obtain the counterfactual assignment of the applicants to programs $\hat{\pi}(X_{i},Q)\in\mathcal{P}$. Second, we construct the AIPW estimate of the “policy value,” that is, the outcome value when treatments are assigned according to $\hat{\pi}(X_{i},Q)\in\mathcal{P}$. We do that using cross-fitting at the applicant level in the test sample, i.e., we estimate new outcome and propensity models and obtain AIPW estimates of the policy values in held-out folds; specifically, we use 100 repetitions with 5 folds and consider the median outcome across repetitions in (chernozhukov2017double).
@@
Throughout this section, we combine the two experimental data sets with data on non-selected applicants to Mentoring and late applicants to Challenges. Thus, we have 997 applicants who did not participate in any program within our sample, resulting in a total sample of 1314 applicants and 19,000 applicant-month observations. For both policy estimation and evaluation, we use survival forests to estimate outcome models and regression forests for the propensity models. The propensity model adjusts for differences between participants and non-participants determined by selection to Mentoring and differences in the populations of applicants to Mentoring and to Challenges. Appendix H provides details of the models used in policy estimation. Appendix I, presents details of the models used in the policy evaluation, the cross-fitting procedure, and validates the approach by comparing the off-policy estimates with the randomized controlled trial estimates. Our procedure recovers the experimental estimates.
@@
In all the preceding analyses, we measured treatment effects as changes in the probability of securing technology employment. For policy evaluation, we adopt an alternative metric that captures additional dimensions of program impact: the restricted mean survival time (RMST), which measures the average time until technology job acquisition over our observation period. The RMST metric offers two advantages for comparing policies. First, it incorporates both whether participants find jobs and how quickly they do so, a policy that helps participants find jobs faster, even if it doesn’t change the ultimate employment probability, generates economic value. Second, RMST estimates aggregate more naturally across heterogeneous populations and policy scenarios, facilitating comparisons of the assignment rules. We report RMST gains in months: negative values indicate participants find technology employment sooner relative to non-participants. An RMST gain of -3 months means treated participants find technology jobs an average of three months earlier. We use survival forests to estimate average treatment effects (athey2019generalized; cui2023estimating).
@@ 6.1 Targeted Assignment Policies
Let $z_{ip}\in\left\{0,1\right\}$ be a variable that indicates whether applicant $i$ is assigned to program $p\in P$, and $z_{ip}^{*}$ be the arguments that maximize the sum of treatment effects given by Equation 2:
$\max_{z_{ip}}\sum_{i=1}^{I}\sum_{p\in\mathcal{P}}z_{ip}\tau_{i}^{p}\text{ s.t. }\sum_{i=1}^{I}z_{ip}\leq Q^{p}\thickspace\forall p\text{ \& }\sum_{p\in\mathcal{P}}z_{ip}=1\thickspace\forall i,$ (2)
where $\tau_{i}^{p}=\mathbb{E}\left[Y(p)-Y(O)\right]$. The first constraint ensures that capacity limits are not violated, and the second is that each applicant is assigned to one allocation. A policy that induces allocation $z_{ip}^{*}$ is what we call the targeted policy. $\tau_{i}^{p}$’s are estimated using the AIPW estimator trained on the train set. In Algorithm 1 in Appendix H, we provide further details of the estimation of the targeted policy.
The outcome of interest, policy value, is the participants’ average treatment effect given program allocation induced by the policy $\pi$. Specifically, we use the AIPW estimator to estimate the average difference in outcomes between being selected to participate based on the assignment policy and not participating in any Dare IT program. This emphasis on impact, rather than average outcomes of participants, aligns with Dare IT’s mission to create the greatest benefit for their participants.
@@ 6.2 Value of Targeted Policies
Given limited program capacity, which applicants should be assigned to which program to maximize overall impact? We estimate the targeted policy at three capacity levels: the 15/15/70 policy approximates existing capacity (assigning 15% of all applicants to Mentoring and 15% to Challenges); the 33/33/34 policy increases capacity to 33% for each program; and the 50/50/0 policy allows each program to admit 50% of applicants, ensuring a program slot for every applicant. We compare these policies against random allocation as a benchmark (random policy allocates all applicants at random to either Mentoring or Challenges). In Section 6.3, we examine how mentor-based selection compares to the algorithm-driven policy in greater detail. The policies vary in both capacity levels and whether participants are systematically targeted.
@@
To understand which applicant types are assigned to each program under data-driven targeting, Figure 3 presents standardized differences in baseline characteristics between optimally assigned individuals and those not assigned to treatment. We implement the 15/15/70 targeted allocation policy and compare covariate distributions of each assigned group to the not-assigned group. Circles represent comparisons for Mentoring; triangles represent Challenges. Dashed vertical lines mark the $\pm$0.1 threshold for imbalance.
Figure 3: Characteristics of Optimally Assigned Groups
@@
Notes: Standardized differences in baseline characteristics between optimally assigned individuals and those not assigned under the 15/15/70 allocation policy. Circles show Mentoring vs Not Assigned; triangles show Challenges vs Not Assigned. The solid vertical line at zero indicates perfect balance; dashed lines mark the $\pm$0.1 threshold. Standardized differences are calculated as the difference in means divided by the pooled standard deviation. Positive values indicate the optimally assigned group has higher values on that characteristic.
@@
The targeted allocation exhibits clear differentiation in targeting patterns across programs. Applicants selected for Mentoring are more likely to hold STEM degrees, live outside Warsaw, have Master’s degrees, and be over 30 years old. Notably, although the algorithm does not use information about prior mentor selection, previously mentor-selected applicants are disproportionately targeted by the algorithm—suggesting the model captures similar signals to those used by mentors. For Challenges, the algorithm shows more balanced selection across characteristics. Assigned applicants are moderately more likely to have been previously selected by mentors but are less likely to hold STEM degrees, live in small towns, or have extensive professional experience compared to those not selected.
Main result.
Table 6 shows gains from data-driven targeting, with effect sizes varying by program and capacity constraint. Under the most selective policy (15/15/70), the overall treatment effect reaches 4.34-4.34 months, which corresponds to an 86% improvement over random allocation’s 2.33-2.33 months. These gains exhibit a strong capacity-selectivity tradeoff: as capacity expands from 15% to 50%, per-person treatment effects decline (from 4.34-4.34 to 2.86-2.86 months overall), though even universal coverage (50/50/0) achieves effects 23% larger than random allocation. Thus, we find that there are gains from targeting across programs.
@@
The targeting gains differ across programs. Mentoring exhibits the largest effects under selective policies (5.55-5.55 months at 15% capacity versus 2.31-2.31 under random), while Challenges shows more modest differentiation (3.13-3.13 versus 2.35-2.35). Notably, even at 50% capacity where the algorithm must accept marginal applicants with lower predicted treatment effects, Mentoring maintains treatment effects 54% larger than random allocation, demonstrating that observable characteristics contain substantial predictive power for treatment response beyond what random assignment captures.
Table 6: Targeted Allocation Policies: Treatment Effects by Program
Capacity Allocation Mentoring Challenges Overall
(Ment. / Chal. / None) Effect Effect Effect
@@
Notes: Treatment effects measured as change in months to employment (negative values = faster job finding). Standard errors in parentheses. Each row shows a capacity-constrained allocation policy specifying the share of applicants assigned to Mentoring, Challenges, or neither program. Random Allocation assigns each program to a random 50% of applicants.
Table 7 examines treatment effects within each assignment group under the 15/15/70 targeted policy. An assignment group consists of all applicants allocated to a specific program (Mentoring, Challenges, or neither) under this policy. For each group, we estimate counterfactual treatment effects under alternative program assignments to assess whether the targeted allocation correctly matches individuals to their highest-benefit program.
The results reveal effective targeting. The Mentoring assignment group receives large benefits from their targeted allocation (5.55-5.55 months, SE = 0.086), substantially exceeding what they would gain from Challenges (3.12-3.12 months, SE = 0.062). The difference of 2.43 months (SE = 0.106) is statistically significant. The Challenges assignment group also benefits most from their assigned program (3.13-3.13 months, SE = 0.069) compared to Mentoring (1.87-1.87 months, SE = 0.050); a difference of 1.26 months (SE = 0.085). This pattern demonstrates benefits from targeting across both programs rather than only selecting to a program.
The Out of Dare IT group (70% of applicants receiving no program assignment) exhibits substantially smaller treatment effects under either counterfactual assignment: 1.72-1.72 months (SE = 0.024) for Mentoring and 2.02-2.02 months (SE = 0.022) for Challenges. While this group would experience gains from program participation, their predicted effects are smaller than those optimally assigned to treatment.
Table 7: Gains From Assignment Policy per Assignment Group
Mentoring group Challenges group Out of Dare IT group
Share 0.15 0.15 0.70
ATE under assignment to:
Targeted -5.552 (0.086) -3.132 (0.069)
Mentoring -5.552 (0.086) -1.872 (0.050) -1.720 (0.024)
Challenges -3.122 (0.062) -3.132 (0.069) -2.017 (0.022)
Difference in ATE between assignments:
Mentoring - Challenges -2.431 (0.106) 1.259 (0.085) 0.297 (0.032)
Note: ATEs are cross-fitted AIPW estimates from causal survival forests (RMST changes up to the evaluation horizon). Negative values indicate earlier job finding (months sooner) relative to out of Dare IT programs. Standard errors in parentheses are GRF mean-of-variance plug-in SEs aggregated within arm.
@@
When comparing the average treatment effects under targeted assignments and the random assignment, we find that there are subsets of applicants who benefit substantially from targeting. However, when the capacity of the two programs increases, the additional participants who get in have lower treatment effects, as suggested by the comparison of targeted policies in Table 6. Thus, there are diminishing returns to targeting additional slots. To identify the level of capacity that realizes most of the available gains from targeting, we estimate the outcomes from targeted assignment under a sequence of capacity values and compare the resulting average probability of getting a job in technology with this outcome under random assignments to Mentoring and to Challenges.
Figure 4: Policy Value Across Capacity Levels: Targeted vs Random Allocation
Notes: Average treatment effects under targeted and random allocation policies as capacity varies from 5% to 50% per program. The solid curve shows targeted allocation that assigns individuals with highest predicted treatment effects to each program; the dashed line shows random assignment of 50% to each program. Treatment effects measured as change in months to employment (negative = faster job finding). Shaded region indicates 95% confidence interval for targeted policy. Policy value evaluated using cross-fitted AIPW estimates on test set. At each capacity level, targeted allocation maximizes expected policy value by solving the assignment problem in Equation 2.
Figure 4 presents the capacity-selectivity tradeoff by plotting average treatment effects as program capacity varies from 5% to 50% per program. The solid curve shows targeted allocation; the dashed horizontal line represents random assignment. Several patterns emerge. First, targeted allocation consistently outperforms random allocation across all capacity levels, with gains largest when capacity is most constrained. At 15% capacity per program, the targeted policy achieves treatment effects of 4.3-4.3 months compared to random allocation’s 2.3-2.3 months, a 2-month improvement. Second, the targeted policy exhibits sharply diminishing returns as capacity expands: per-person effects decline from approximately 4.5-4.5 months at 5% capacity to 2.9-2.9 months at 50% capacity. This reflects the algorithm’s sequential assignment of highest-benefit individuals first, with marginal participants exhibiting progressively smaller treatment effects. Third, even at universal coverage (50% per program, accommodating all applicants), targeted allocation maintains a 0.6-month advantage over random assignment—demonstrating that complementarity between individual characteristics and program type generates value beyond pure capacity expansion.

### v5

@@ 6 Counterfactual Program Assignment Policies
Program capacity is limited. Mentor availability restricts Mentoring capacity, while we capped Challenges at 300 participants to manage operational complexity. Given the substantial treatment effects documented above, relaxing these constraints would increase Dare IT’s impact. However, capacity constraints are likely to persist, particularly for Mentoring. Dare IT therefore faces an ongoing allocation problem: which applicants should be admitted to each program? The heterogeneous treatment effects documented in Section 5.3 suggest that admission policies favoring applicants with high predicted treatment effects could enhance program impact. This section quantifies potential gains from such targeting, characterizes which applicant profiles benefit most from each program, and examines whether the patterns align with the theoretical mechanisms identified above.
@@
We perform three analyses: (i) we estimate gains from an assignment policy that allocates applicants to programs to maximize the sum of predicted treatment effects, given capacity constraints; (ii) we evaluate gains from relaxing capacity constraints with this targeting policy in place; and (iii) we compare overall Mentoring impact under mentor-based selection to alternative assignment strategies.
Policy framework.
@@
An assignment policy maps applicant characteristics $X\in\mathcal{X}$ and program capacities $Q=(Q^{M},Q^{C})\in\mathcal{Q}$ to assignments: $\pi:\mathcal{X}\times\mathcal{Q}\rightarrow\mathcal{P}$, where $\mathcal{P}\equiv\{M,C,O\}$ denotes Mentoring, Challenges, and Out of Dare IT, respectively. $\mathcal{Q}$ is the grid of capacity levels under consideration.
@@
We perform policy estimation and evaluation in separate samples. We randomly split the data at the applicant level with equal probability into training and test sets ($\mathcal{I}_{\text{train}}$ and $\mathcal{I}_{\text{test}}$). Using the training set, we estimate the policy $\hat{\pi}$. To evaluate the policy, we apply $\hat{\pi}$ to the test set, obtaining counterfactual assignments $\hat{\pi}(X_{i},Q)\in\mathcal{P}$ for each applicant $i$ and capacity level $Q$. We then construct augmented inverse propensity weighted (AIPW) estimates of the policy value—the expected outcome when treatments are assigned according to $\hat{\pi}$. We implement this using cross-fitting at the applicant level in the test sample: we estimate outcome and propensity models and obtain AIPW estimates in held-out folds, using 100 repetitions with 5 folds and reporting the median across repetitions (Chernozhukov et al., 2017).
Sample and estimation.
@@
We combine the two experimental samples with non-selected applicants to Mentoring and late applicants to Challenges, yielding 1,314 total applicants (682 in experimental samples, 632 non-participants) and 19,000 applicant-month observations. We use survival forests to estimate outcome models and regression forests for propensity models. The propensity model adjusts for selection into Mentoring and for differences between the Mentoring and Challenges applicant populations. Appendix J details the estimation procedure; Appendix K describes the evaluation models and cross-fitting procedure, and validates the approach by comparing off-policy estimates to experimental estimates.
@@
The preceding analyses measured treatment effects as changes in the probability of technology employment. For policy evaluation, we adopt the restricted mean survival time (RMST), which measures the average time until technology job acquisition over our observation period. RMST incorporates both whether participants find jobs and how quickly they do so: a policy that accelerates job-finding generates economic value even if it does not change ultimate employment probability.
We report RMST effects in months, with negative values indicating faster job acquisition. An RMST effect of 3-3 months means treated participants find technology jobs three months earlier on average than non-participants. We estimate treatment effects using causal survival forests (Athey et al., 2019; Cui et al., 2023). Appendix I presents average treatment effect estimates in the experimental samples using this method.
@@ 6.1 Targeted Assignment Policies
We formulate optimal program assignment as an integer linear program. Let $z_{ip}\in\{0,1\}$ indicate whether applicant $i$ is assigned to program $p\in\mathcal{P}$, where $\mathcal{P}=\{M,C,O\}$ denotes Mentoring, Challenges, and no program (Out), respectively. The optimal assignment solves:
$\max_{z_{ip}}\sum_{i=1}^{N}\sum_{p\in\mathcal{P}}z_{ip}\tau_{i}^{p}\quad\text{s.t.}\quad\sum_{i=1}^{N}z_{ip}\leq Q^{p}\;\forall p,\quad\sum_{p\in\mathcal{P}}z_{ip}=1\;\forall i,$ (3)
where $\tau_{i}^{p}=\mathbb{E}[Y_{i}(p)-Y_{i}(O)]$ is applicant $i$’s treatment effect from program $p$ relative to no program, and $Q^{p}$ is the capacity of program $p$ (with $Q^{O}=\infty$). The first constraint ensures capacity limits are respected; the second ensures each applicant receives exactly one assignment.
We call the policy inducing the solution $z_{ip}^{*}$ the targeted policy. Individual treatment effects $\tau_{i}^{p}$ are estimated using AIPW with models trained on the training set. Algorithm 1 in Appendix J provides implementation details.
To evaluate policies, we estimate the policy value: the average treatment effect achieved when applicants are assigned according to policy $\pi$ rather than receiving no program. This focus on program impact—the causal effect of participation—rather than raw employment rates among participants aligns with Dare IT’s goal of maximizing benefit to participants. Below, we compare targeting based on predicted treatment effects to alternative approaches: selecting applicants with the lowest baseline employment probability or those with the highest baseline probability.
@@ 6.2 Value of Targeted Policies
Given limited program capacity, which applicants should be assigned to which program to maximize overall impact? We estimate targeted policies at three capacity levels, where capacity is expressed as the percentage of applicants assigned to Mentoring/Challenges/Out (no program): 15/15/70 approximates existing capacity; 33/33/34 expands both programs moderately; and 50/50/0 ensures every applicant receives a program slot. As a benchmark, we compare these targeted policies against random assignment, which allocates applicants to Mentoring or Challenges with equal probability. We evaluate the random policy at full capacity (50/50/0) because lower capacity levels yield the same expected treatment effects but noisier estimates due to fewer treated subjects.
@@
To understand which applicant types the algorithm assigns to each program, Figure 3 presents standardized differences in baseline characteristics between optimally assigned applicants and those assigned to no program under the 15/15/70 policy. Circles represent Mentoring; triangles represent Challenges. Dashed vertical lines mark the $\pm 0.1$ threshold conventionally used to indicate meaningful imbalance.
Figure 3: Characteristics of Groups Under Targeted Policy
@@
Notes: Standardized differences in baseline characteristics between applicants assigned to each program and those assigned to no program under the 15/15/70 targeted policy. Circles show Mentoring versus Out; triangles show Challenges versus Out. Dashed lines mark the $\pm 0.1$ threshold. Standardized differences are calculated as the difference in means divided by the pooled standard deviation; positive values indicate the assigned group has higher values on that characteristic. Selected by Mentor is not used in the policy assignment, presented here for comparison.
@@
The algorithm exhibits clear differentiation in targeting across programs. Applicants assigned to Mentoring are more likely to hold STEM degrees, live outside Warsaw, have Master’s degrees, and be over 30 years old. These patterns align with the heterogeneous treatment effects documented in Section 5.3, where Mentoring showed larger effects for participants outside Warsaw and those without STEM backgrounds. Although the algorithm does not observe prior mentor selection status, applicants whom mentors selected in the experiment are disproportionately assigned to Mentoring, suggesting the algorithm captures similar signals to those mentors used when choosing mentees.
For Challenges, targeting is more balanced across characteristics. Assigned applicants are moderately more likely to have been mentor-selected but are less likely to hold STEM degrees, live in small towns, or have extensive professional experience. This pattern is consistent with Challenges benefiting early-career women who lack credentials to signal their skills, as documented in the HTE analysis.
Main results.
Table 6 presents treatment effects under data-driven targeting at varying capacity levels, estimated over the pooled population of experimental participants from both programs and all other eligible applicants (non-selected Mentoring applicants and late Challenges applicants). Recall that negative values indicate faster job finding (months saved). Under the most selective policy (15/15/70), the overall treatment effect reaches 4.34-4.34 months, an 86% improvement over random allocation’s 2.33-2.33 months. These gains exhibit a capacity-selectivity tradeoff: as capacity expands from 15% to 50% per program, average treatment effects decline from 4.34-4.34 to 2.86-2.86 months as the algorithm must include applicants with lower predicted benefits. Even at universal coverage (50/50/0), however, targeted assignment achieves effects 23% larger than random allocation. [^10] Appendix L presents an alternative specification using repeated cross-fitting following Fava (2025), which uses all observations for both policy estimation and evaluation rather than separate training and test samples. Point estimates are nearly identical (e.g., 4.60-4.60 versus 4.34-4.34 months for the 15/15/70 policy), while standard errors are smaller due to efficiency gains from the larger effective sample size. The qualitative conclusions are unchanged.
@@
Targeting gains differ substantially across programs. Mentoring exhibits the largest effects under selective policies: 5.55-5.55 months at 15% capacity versus 2.31-2.31 under random assignment. Challenges shows more modest gains from targeting (3.13-3.13 versus 2.35-2.35 months, a 33% improvement). Even at 50% capacity, targeted allocation achieves Mentoring treatment effects of -3.55 months compared to -2.31 months under random allocation (54% larger), demonstrating that observable characteristics contain substantial predictive power for treatment response.
Table 6: Treatment Effects Under Targeted Allocation Policies
Mentoring Challenges Overall
@@
Notes: Treatment effects measured as change in restricted mean survival time (months); negative values indicate faster job finding. Capacity allocations specify the share of applicants assigned to Mentoring/Challenges/Out. Random allocation assigns each applicant to either program with equal probability. Standard errors in parentheses.
Table 7 examines whether targeting correctly matches applicants to their highest-benefit program. We estimate counterfactual treatment effects for each assignment group under the 15/15/70 policy. In other words, we estimate, what would applicants assigned to Mentoring gain if instead assigned to Challenges, and vice versa?
The results confirm effective assignment. Applicants assigned to Mentoring gain 5.55-5.55 months from their assignment but would gain only 3.12-3.12 months from Challenges—a difference of 2.43 months (SE 0.11). Applicants assigned to Challenges similarly benefit more from their assigned program (3.13-3.13 months) than from Mentoring (1.87-1.87 months), a difference of 1.26 months (SE 0.09). The algorithm thus identifies not only who benefits most from treatment, but which treatment benefits each applicant most.
Applicants assigned to no program (70% of the pool) exhibit smaller predicted effects under either counterfactual: 1.72-1.72 months for Mentoring and 2.02-2.02 months for Challenges. While these applicants would benefit from participation, their predicted gains are smaller than those of optimally assigned applicants. Interestingly, this group would benefit relatively more from Challenges than Mentoring, suggesting that among lower-priority applicants, the lower-cost program delivers comparable or greater value.
Table 7: Counterfactual Treatment Effects by Assignment Group
Assigned to Mentoring Assigned to Challenges Assigned to Out
Share of applicants 0.15 0.15 0.70
Treatment effect if assigned to:
Mentoring -5.55 (0.09) -1.87 (0.05) -1.72 (0.02)
Challenges -3.12 (0.06) -3.13 (0.07) -2.02 (0.02)
Difference:
Mentoring - Challenges -2.43 (0.11) 1.26 (0.09) 0.30 (0.03)
Notes: Treatment effects measured as change in restricted mean survival time (months); negative values indicate faster job finding. Each column represents applicants assigned to that program under the 15/15/70 targeted policy. Rows show estimated treatment effects if that group were counterfactually assigned to each program. Standard errors in parentheses.
@@
The preceding results examine discrete capacity levels. To characterize the full capacity-selectivity tradeoff, Figure 4 plots average treatment effects as program capacity varies continuously from 5% to 50% per program. The solid curve shows targeted allocation; the dashed horizontal line represents random assignment.
Figure 4: Policy Value Across Capacity Levels
Notes: Average treatment effects (RMST, in months) under targeted allocation as capacity varies from 5% to 50% per program. Shaded region indicates 95% confidence interval. Policy value evaluated using cross-fitted AIPW estimates on the test set. At each capacity level, targeted allocation solves the assignment problem in Equation 3. Dashed line shows the average treatment effect under random assignment with equal probability across programs.
Three patterns emerge. First, targeted allocation outperforms random assignment at every capacity level, with the largest gains when capacity is most constrained. At 5% capacity, targeting achieves effects of approximately 4.5-4.5 months compared to random allocation’s 2.3-2.3 months. Second, returns to targeting diminish as capacity expands: per-person effects decline steadily as the algorithm exhausts high-benefit applicants and must include those with smaller predicted gains. Third, and most notably, targeted allocation maintains a meaningful advantage even at universal coverage (50% per program). The 0.6-month improvement at full capacity reflects gains from matching applicants to their best-suited program, not just from selecting high-benefit individuals. This complementarity between applicant characteristics and program type generates value independent of capacity constraints.